A MAN IS 27 YEARS OLDER THAN HIS SON , AND IN 10 YEARS FROM NOW HE WILL BE TWICE AS OLD AS HIS SON . HOW OLD IS EACH NOW ?

Answer by Maths68(1474)

Let
Present age of Man = m years old
Present age of Son = s years old

Man is 27 years older than his son
m=s+27...........(1)

In 10 years
Man will be = m+10 years old
Son will be = s+10 years old
Then
Man will be twice as old as his Son
m+10=2(s+10)
m+10=2s+20
m=2s+20-10
m=2s+10...............(2)
Substitute the value of m from (1) to (2)
s+27=2s+10
s-2s=10-27
-s=-17
Multiply by -1 both sides
s=17
Substitute the value of s in (1)
m=s+27
m=17+27
m=44

Present age of Man = m = 44 years old
Present age of Son = s = 17 years old